OpenStudy (anonymous):

Solve for X. x-2/3=2 2/3

OpenStudy (anonymous):

is that 8/3 or 22/3?

OpenStudy (anonymous):

2 2/3

OpenStudy (anonymous):

2 and two over three

OpenStudy (anonymous):

ok, so then \[x=10/3\] because you make it so its \[x-2/3=8/3\] and you add 2/3 to both sides

OpenStudy (anonymous):

no its not x-3/4=2 2/4

OpenStudy (anonymous):

sorry its x-3/4=2 2/3

OpenStudy (anonymous):

\[x-3/4=8/3\] make demoninators equal:\[x-9/12=32/12\] \[x=41/12\]

OpenStudy (anonymous):

its 25/12

OpenStudy (anonymous):

are you sure the problem is the right problem?